OpenStudy (anonymous):

what is the solution to the proportion 9/x=4/12

OpenStudy (qwertty123):

Do you know how to start?

OpenStudy (anonymous):

no help

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

Done qwerty

OpenStudy (anonymous):

can u help?

OpenStudy (qwertty123):

yes

OpenStudy (qwertty123):

9*12=108 4*x=4x so, 4x/108 4x=108 x=?

OpenStudy (anonymous):

qwert do me

OpenStudy (anonymous):

idk qwert

OpenStudy (anonymous):

Shes right

OpenStudy (anonymous):

ik but like what is x

OpenStudy (qwertty123):

do 108/4!

OpenStudy (anonymous):

i got it thank you it what is 27 follow me on instagram now htx_dominic1023

OpenStudy (qwertty123):

You welcome!

OpenStudy (cmtboy2):

Dominic is my name

OpenStudy (anonymous):

samee

OpenStudy (qwertty123):

:)

OpenStudy (anonymous):

follow me

OpenStudy (cmtboy2):

my full name is Dominic russell freeman

OpenStudy (qwertty123):

dude I don't have that

OpenStudy (anonymous):

qwert cmtboy go to first question

OpenStudy (qwertty123):

okay

OpenStudy (anonymous):

13/20 as a decimal

OpenStudy (anonymous):

??

OpenStudy (anonymous):

nvm its .65

OpenStudy (anonymous):

A bag contains 12 red marbles and 4 blue marbles. What is the ratio, written in lowest terms, of red marbles to the total number of marbles in the bag?